# Reinforced MLLM: A Survey on RL-Based Reasoning in Multimodal Large Language Models

### v1

@@ 6. Datasets and Benchmark
Insight-V 59.9 - - 50.2 61.5 1685.1 81.5 - - -
LLaVA-Reasoner 50.6 - - - - - 83.0 - - -
Kimi K1.5-Long CoT 74.90 38.60 - 70.00 - - - - - -
Kimi K1.5-Short CoT 70.10 31.00 - 68.00 - - - - - -
Vision-R1-7B 73.50 - 52.40 - - - - - - -
Vision-R1-LlamaV-CI-11B 62.70 - 27.10 - 61.40 2190.00 83.90 - 49.50 -
MM-Eureka-8B 67.10 22.20 40.40 - - - - - - 8.60
Curr-ReFT-3B 58.60 - - - - - - - - -
Curr-ReFT-7B 64.50 - - - - - - - - -
LMM-R1: MGT-Multi 59.00 26.84 41.80 - 54.39 - - - - 14.63
LMM-R1: MGT-PerceReson 63.20 26.35 41.55 - 58.03 - - - - 15.62
R1-Onevision-7B 64.10 29.90 40.00 - - - - - - -
R1-VL-2B 52.10 17.10 26.20 - 49.80 2048.00 75.20 29.40 44.00 -
R1-VL-7B 63.50 24.70 40.00 - 60.00 2376.00 83.90 45.20 54.70 -
Skywork R1V 67.50 - - 69.00 - - - - - -
@@ 6.1. Reasoning over Structured Knowledge Domains
Chart-based reasoning benchmarks evaluate the ability of MLLMs to extract structured information from visual plots and perform quantitative analysis grounded in natural language. ChartBench (Xu et al., 2023) and MultiChartQA (Zhu et al., 2024) present a wide range of questions over diverse chart types, such as bar graphs, pie charts, and line plots, requiring aggregation, comparison, or prediction over visual data. CharXiv (Wang et al., 2024c) adopts a more challenging evaluation setting by sourcing all charts from real-world scientific publications. The benchmark incorporates complex domain-specific notations, multi-subplot layouts, and visually dense figures, as well as semantically ambiguous or unanswerable questions, thereby substantially increasing the task complexity. This dataset collectively assesses the model’s comprehensive capabilities in multimodal alignment, numerical reasoning, and visual analytical inference.
@@
Scientific reasoning benchmarks aim to assess a model’s ability to integrate domain-specific visual information with conceptual understanding across disciplines such as physics, chemistry, and biology. MMMU (Yue et al., 2024a) and MMMU-Pro (Yue et al., 2024b) provide large-scale multimodal assessments based on academic exam questions, with inputs that include circuit diagrams, biological illustrations, and chemical reaction graphs. These benchmarks demand subject-specific knowledge and cross-modal multi-hop inference. M4U (Wang et al., 2024d) extends this direction by constructing a multilingual, multi-discipline benchmark that includes 64 subjects and supports both English and Chinese. Its tasks feature rich visual materials such as graphs, experimental setups, and symbolic content, paired with multilingual questions designed to test deep scientific understanding, multilingual alignment, and multimodal reasoning in real-world educational settings. ScienceQA (Lu et al., 2022) and TheoremQA (Chen et al., 2023a) offer grade-school to high-school level science problems that emphasize explanation and factual integration, often including illustrated contexts or structured diagrams. GAOKAO-MM (Zong and Qiu, 2024) and CMMMU (Ge et al., 2024) reflect the format of high-stakes standardized tests, containing visually dense problem structures with minimal textual scaffolding. EMMA (Hao et al., 2025) further enhances the realism of scientific instruction by combining textual questions with experimental setups or process diagrams, requiring temporal reasoning and procedural interpretation. To address the issue of visual leakage in multimodal evaluation datasets, MM-Star (Chen et al., 2024b) assesses the multimodal capabilities of LVLM by constructing a carefully balanced and purified sample selection process, with the aim of benchmarking six core competencies and 18 detailed axes.
@@
Benchmarks in this category evaluate the ability of MLLMs to generate code based on multimodal visual prompts, such as plots or interface sketches. ChartMimic (Shi et al., 2024) and Plot2Code (Wu et al., 2024b) require models to produce plotting scripts (e.g., matplotlib) that replicate the semantics and structure of a given graph. These tasks test visual layout parsing, semantic inference, and parameter mapping between modalities. HumanEval-V (Zhang et al., 2024c) extends this framework by offering open-ended coding tasks guided by visual problem descriptions, where models must synthesize intent, logic, and implementation through image-grounded reasoning.
@@
The NPHardEval4V (Fan et al., 2024) benchmark focuses on visually rendered algorithmic reasoning tasks, including classical problems such as graph coloring, knapsack optimization, and shortest path computation. The tasks require interpreting structured visual representations—such as graphs, grids, or symbolic layouts—and solving them via rule-based reasoning and efficient search. This setting probes a model’s ability to map visual configurations to algorithmic abstractions and reason over computationally complex solution spaces.
@@
This subcategory evaluates MLLMs’ ability to reason about sociocultural, geographic, and symbolic content presented in visual form. II-Bench (Liu et al., 2024a), CII-Bench (Zhang et al., 2024a), PunchBench (Ouyang et al., 2024), and GeoComp (Song et al., 2025) introduce tasks involving photographs, illustrations, posters, and maps, where models must interpret not only the visual content but also its embedded cultural or societal implications. CII-Bench (Zhang et al., 2024a) features Chinese cultural imagery including traditional art and contemporary symbolism, probing metaphor understanding and emotional alignment. PunchBench (Ouyang et al., 2024) and GeoComp (Song et al., 2025) emphasize spatial-social reasoning grounded in world knowledge and demographic cues. These datasets collectively assess whether models can move beyond surface-level perception to perform culturally informed and context-sensitive inference.
@@ 6.2. Contextual and Perceptual Reasoning in Real-World Scenarios
Benchmarks in this subcategory simulate scenarios where agents must make predictions or decisions based on dynamic visual contexts. Driving-Hazard-Prediction-and-Reasoning (Charoenpitaks et al., 2024) presents real-world traffic images and sequential scenarios to assess hazard identification and behavioral prediction. InfiGUIAgent-Data (Liu et al., 2025b), Mind2Web (Deng et al., 2023), and SeeClick (Cheng et al., 2024) extend this to user-interface environments, requiring models to reason about user intent and anticipate interactions with graphical user interfaces. Mind2Web (Deng et al., 2023) further introduces a broad set of high-level web-based tasks grounded in real-world websites, evaluating models on instruction following, multi-step planning, and interface understanding. Collectively, these benchmarks examine MLLMs’ capacity to reason about actions and decisions within temporally evolving or interactive settings.
@@
Spatial reasoning benchmarks target the model’s understanding of object layout, spatial relationships, and three-dimensional scene structure. iVISPAR (Mayer et al., 2025) evaluates MLLMs on spatial visual question answering tasks involving object localization, spatial relation inference, and layout understanding in cluttered indoor scenes. It emphasizes relational reasoning and occlusion modeling based on realistic 2D imagery. PulseCheck457 (Wang et al., 2025b) advances this line of work by introducing the first benchmark explicitly designed for 6D spatial reasoning. It assesses models on multi-object recognition, 2D and 3D positioning, and 3D orientation within a synthetic, unbiased environment. Tasks span five difficulty levels and seven question types—including occlusion detection, orientation alignment, and collision prediction—offering structured diagnosis of MLLMs’ spatial abstraction abilities for embodied AI, robotics, and AR/VR applications.
@@ 6.3. Generalization and Multi-Task Reasoning Benchmarks
This category includes composite benchmarks designed to assess MLLMs’ abilities in multi-task reasoning, cross-domain generalization, and zero-shot adaptation. These datasets typically aggregate diverse task types across multiple modalities, testing whether models can perform structured inference under minimal supervision. Tasks range from multimodal CoT generation to inductive abstraction and analogical reasoning, providing a comprehensive evaluation of models’ capacity to generalize beyond domain-specific constraints and execute flexible reasoning strategies across heterogeneous inputs.
@@
To evaluate inductive abstraction and rule discovery, benchmarks like MM-IQ (Cai et al., 2025), LogicVista (Xiao et al., 2024), and PuzzleTest (Toh et al., 2025) present visual matrix-style puzzles inspired by Raven’s Progressive Matrices. These datasets challenge models to infer abstract transformations, detect visual analogies, and fill in missing patterns by synthesizing structural invariants across multiple image frames. They emphasize high-level visual abstraction and the ability to generalize across non-linguistic reasoning patterns.
@@
A growing number of composite benchmarks have been proposed to evaluate generalized multimodal reasoning across diverse task types and domains. Representative datasets such as VRC-Bench (Thawakar et al., 2025) and MLLM-COMPBench (Kil et al., 2024) aggregate subtasks including visual question answering, multi-hop reasoning, and tool-augmented inference, providing a unified setting for multi-skill evaluation. Notably, M3CoT (Chen et al., 2024d) and MME-CoT (Jiang et al., 2025) incorporate multimodal chain-of-thought annotations, enabling fine-grained process supervision throughout inference. Meanwhile, ZeroBench (Roberts et al., [n. d.]) focuses on zero-shot generalization across unseen tasks and modalities, assessing the model’s transferability and robustness without explicit instruction tuning. Together, these benchmarks form a comprehensive suite for evaluating holistic, structured, and cross-modal reasoning abilities under diverse and challenging input conditions.
@@ 7.1. Limitations and Challenges
One of the most pressing issues is the sparsity and delayed nature of reward signals. To avoid reward hacking, most RL-enhanced MLLMs are trained using scalar final-task rewards—such as answer correctness or classification accuracy—that only reflect task outcomes but offer no intermediate feedback throughout the reasoning process. This leaves the model blind to structural errors in earlier reasoning steps. The most immediate consequence of this issue is overthinking, where the model generates excessively long or redundant reasoning paths, with intermediate steps potentially including irrelevant visual or textual cues. Although recent works like R1-VL and Curr-ReFT have introduced process-oriented feedback through intermediate step verification or curriculum-style stage partitioning, these designs are often handcrafted, task-specific, and difficult to generalize across unseen modalities or new domains.
@@
Another limitation lies in the benchmark-centric and static nature of current evaluation paradigms. Many RL-based MLLMs are trained and tested on a limited range of benchmarks such ScienceQA and MathVista. These datasets are typically narrow in scope and fail to reflect the complexity and unpredictability of real-world multimodal reasoning. As a result, models trained on such benchmarks often exhibit poor transferability when applied to dynamic environments or when new modalities such as audio or 3D spatial layouts are introduced. This issue is particularly pronounced for small-scale models, where limited parameter capacity amplifies the brittleness of reward optimization and modality fusion.
@@
In addition to structural limitations, MLLM RL pipelines face several pressing challenges spanning optimization theory, multimodal alignment, and training infrastructure. One major issue is achieving effective cross-modal alignment under weak supervision. Many real-world tasks require complex mappings between modalities such as images, text, audio, and spatial layouts, yet direct supervision for these transformations is often limited. Designing reward functions that can consistently align such diverse modalities—particularly in open-ended or underspecified scenarios—remains a significant open problem. Another challenge lies in handling non-Markovian dependencies within reasoning trajectories. Unlike traditional RL environments, MLLM-based reasoning demands long-term consistency across multiple modalities and sequential reasoning steps. Although frameworks like GFlowVLM and MetaSpatial attempt to model these dependencies, optimizing policies in such non-Markovian spaces leads to instability, increased gradient noise, and difficult credit assignment. Furthermore, there is a notable mismatch between training and inference phases, where models are often trained with fixed prompts and deterministic rewards but evaluated under conditions involving unpredictable inputs, variable-length reasoning chains, and ambiguous outputs. This discrepancy undermines real-world performance despite strong benchmark results. Lastly, scalability remains a bottleneck, as many process-aligned RL methods demand substantial compute for multi-step decoding, intermediate scoring, and CoT supervision. These demands hinder practical deployment, particularly in academic or resource-constrained settings.
@@ 7.2. Future Directions
To overcome these limitations and challenges, we propose several concrete directions for future research that can contribute to more robust, generalizable, and interactive reinforcement learning in MLLMs:
@@
Future work should focus on designing multi-level reward signals that not only reflect task outcomes but also capture the fidelity of the reasoning process and structural coherence across modalities. For example, hybrid reward structures can integrate accuracy (e.g., correctness of final answers), structure (e.g., step order, logical flow), and quality (e.g., aesthetics in image generation, coherence in story generation). Such reward hierarchies can enhance sample efficiency, interpretability, and policy stability.
@@
A promising research direction is the development of lightweight RL techniques suitable for resource-constrained models. This includes curriculum learning schedules, KL-regularized off-policy methods, and contrastive reward estimation. Such strategies can facilitate broader adoption in applied scenarios where deploying large models is infeasible.
@@
Deploying MLLMs in spatially grounded environments (e.g., robotics, AR/VR) requires RL strategies that handle physical constraints, causality, and temporal dynamics. This includes integrating spatial consistency checks, collision detection, and user affordance modeling into the reward computation pipeline, as seen in MetaSpatial.

### v2

@@ 6. Datasets and Benchmark
Insight-V 59.9 - - 50.2 61.5 1685.1 81.50 - - -
LLaVA-Reasoner 50.6 - - - - - 83.00 - - -
Kimi K1.5-Long CoT 74.9 38.6 - 70.0 - - - - - -
Kimi K1.5-Short CoT 70.1 31.0 - 68.0 - - - - - -
Kimi-VL-A3B 68.7 21.4 - 57.0 61.3 - - - - 66.7
Vision-R1-7B 73.5 - 52.4 - - - - - - -
Vision-R1-LlamaV-CI-11B 62.7 - 27.1 - 61.4 2190.0 83.9 - 49.5 -
MM-Eureka-8B 67.1 22.2 40.4 - - - - - - -
Curr-ReFT-3B 58.6 - - - - - - - - 29.9
Curr-ReFT-7B 64.5 - - - - - - - - 35.6
LMM-R1: MGT-Multi 59.0 26.8 41.8 - 54.4 - - - - -
LMM-R1: MGT-PerceReson 63.2 26.4 41.6 - 58.03 - - - - -
R1-Onevision-7B 64.1 29.9 40.0 - - - - - - -
R1-VL-2B 52.1 17.1 26.2 - 49.8 2048.0 75.2 29.4 44.0 -
R1-VL-7B 63.5 24.7 40.0 - 60.0 2376.0 83.9 45.2 54.7 -
Skywork R1V 67.5 - - 69.0 - - - - - -
ThinkLite-VL-7B 75.1 32.9 50.7 54.6 65.00 - - - - 67.8
VLAA-Thinker-Qwen2.5-3B 61.0 24.4 36.4 - - - - - - -
VLAA-Thinker-Qwen2.5-7B 68.0 26.4 48.2 - - - - - - -
VL-Rethinker-7B 74.9 32.3 54.2 56.7 - - - - - -
VL-Rethinker-72B 80.3 43.9 61.7 68.8 - - - - - -
Skywork R1V2 74.0 49.0 - 73.6 - - - - - -
FAST-3B 66.2 26.8 43.0 - - - - 54.4 - 64.0
FAST-7B 73.8 30.6 50.6 - - - - 58.3 - 71.2
NoisyRollout 69.6 28.5 53.2 - - - - - - -
@@ 6.1. Reasoning over Structured Knowledge Domains
Chart-based reasoning benchmarks evaluate the ability of MLLMs to extract structured information from visual plots and perform quantitative analysis grounded in natural language. ChartBench (Xu et al., 2023) and MultiChartQA (Zhu et al., 2024) present a wide range of questions over diverse chart types, such as bar graphs, pie charts, and line plots, requiring aggregation, comparison, or prediction over visual data. CharXiv (Wang et al., 2024c) adopts a more challenging evaluation setting by sourcing all charts from real-world scientific publications. The benchmark incorporates complex domain-specific notations, multi-subplot layouts, and visually dense figures, as well as semantically ambiguous or unanswerable questions, thereby substantially increasing the task complexity.
@@
Scientific reasoning benchmarks aim to assess a model’s ability to integrate domain-specific visual information with conceptual understanding across disciplines such as physics, chemistry, and biology. MMMU (Yue et al., 2024a) and MMMU-Pro (Yue et al., 2024b) provide large-scale multimodal assessments based on academic exam questions, with inputs that include circuit diagrams, biological illustrations, and chemical reaction graphs. These benchmarks demand subject-specific knowledge and cross-modal multi-hop inference. M4U (Wang et al., 2024d) extends this direction by constructing a a multilingual benchmark that covers 64 disciplines. Its tasks feature rich visual materials such as graphs, experimental setups, and symbolic content, paired with multilingual questions designed to test the deep scientific understanding, multilingual alignment, and multimodal reasoning capabilities of MLLMs in real-world educational settings. MDK12-Bench (Zhou et al., 2025b) complements these efforts by focusing on structured academic reasoning across core school subjects, while incorporating dynamic evaluation to enhance robustness. ScienceQA (Lu et al., 2022) and TheoremQA (Chen et al., 2023) offer grade-school to high-school level science problems that emphasize explanation and factual integration, often including illustrated contexts or structured diagrams. GAOKAO-MM (Zong and Qiu, 2024) and CMMMU (Ge et al., 2024) reflect the format of high-stakes standardized tests, containing visually dense problem structures with minimal textual scaffolding. EMMA (Hao et al., 2025) further enhances the realism of scientific instruction by combining textual questions with experimental setups or process diagrams, requiring temporal reasoning and procedural interpretation. To address the issue of visual leakage in multimodal evaluation datasets, MM-Star (Chen et al., 2024a) assesses the multimodal capabilities of LVLM by constructing a carefully balanced and purified sample selection process, with the aim of benchmarking six core competencies and 18 detailed axes.
@@
Benchmarks in this category evaluate the ability of MLLMs to generate code based on multimodal visual prompts, such as plots or interface sketches. ChartMimic (Shi et al., 2024) and Plot2Code (Wu et al., 2024) require models to produce plotting scripts such as matplotlib that replicate the semantics and structure of a given graph. These tasks test visual layout parsing, semantic inference, and parameter mapping between modalities. HumanEval-V (Zhang et al., 2024c) extends this framework by offering open-ended coding tasks guided by visual problem descriptions, where models must synthesize intent, logic, and implementation through image-grounded reasoning.
@@
The NPHardEval4V (Fan et al., 2024) benchmark focuses on visually rendered algorithmic reasoning tasks, including classical problems such as graph coloring, knapsack optimization, and shortest path computation. The tasks require interpreting structured visual representations and aim to explore the model’s ability to map visual configurations to algorithmic abstractions and perform reasoning within computationally complex solution spaces.
@@
This subcategory evaluates MLLMs’ ability to reason about sociocultural, geographic, and symbolic content presented in visual form. II-Bench (Liu et al., 2024a), CII-Bench (Zhang et al., 2024a), PunchBench (Ouyang et al., 2024), and GeoComp (Song et al., 2025) introduce tasks involving photographs, illustrations, posters, and maps, where models must interpret not only the visual content but also its embedded cultural or societal implications. CII-Bench (Zhang et al., 2024a) features Chinese cultural imagery including traditional art and contemporary symbolism, probing metaphor understanding and emotional alignment. PunchBench (Ouyang et al., 2024) and GeoComp (Song et al., 2025) emphasize spatial-social reasoning grounded in world knowledge and demographic cues. These datasets evaluate whether models can go beyond surface-level perception to perform culturally and contextually sensitive reasoning.
@@ 6.2. Contextual and Perceptual Reasoning in Real-World Scenarios
Benchmarks in this subcategory simulate scenarios where agents must make predictions or decisions based on dynamic visual contexts. Driving-Hazard-Prediction-and-Reasoning (Charoenpitaks et al., 2024) presents real-world traffic images and sequential scenarios to assess hazard identification and behavioral prediction. InfiGUIAgent-Data (Liu et al., 2025b), Mind2Web (Deng et al., 2023), and SeeClick (Cheng et al., 2024) extend this to user-interface environments, requiring models to reason about user intent and anticipate interactions with graphical user interfaces. Mind2Web (Deng et al., 2023) further introduces a broad set of high-level web-based tasks grounded in real-world websites, evaluating models on instruction following, multi-step planning, and interface understanding.
@@
Spatial reasoning benchmarks target the model’s understanding of object layout, spatial relationships, and three-dimensional scene structure. iVISPAR (Mayer et al., 2025) evaluates MLLMs on spatial visual question answering tasks involving object localization, spatial relation inference, and layout understanding in cluttered indoor scenes. It emphasizes relational reasoning and occlusion modeling based on realistic 2D imagery. PulseCheck457 (Wang et al., 2025b) advances this work by introducing the first benchmark explicitly designed for 6D spatial reasoning. It assesses models on multi-object recognition, 2D and 3D positioning, and 3D orientation within a synthetic, unbiased environment, covering five difficulty levels and seven question types. This provides structured diagnostics for spatial abstraction capabilities in embodied robotics and AR/VR applications. GeoSense (Xu et al., 2025c) further emphasizes symbolic-visual integration by evaluating geometric principle identification and application in diagram-based reasoning. VisuLogic (Xu et al., 2025b) targets core visual reasoning skills, focusing on symmetry, rotation, folding, and positional changes in structured visual puzzles.
@@ 6.3. Generalization and Multi-Task Reasoning Benchmarks
This category includes composite benchmarks designed to assess MLLMs’ abilities in multi-task reasoning, cross-domain generalization, and zero-shot adaptation. These datasets typically aggregate diverse tasks across modalities, testing whether models can perform structured inference with minimal supervision. Tasks range from multimodal CoT generation to inductive abstraction and analogical reasoning, providing a comprehensive evaluation of models’ capacity to generalize beyond domain-specific constraints and perform flexible reasoning over heterogeneous inputs.
@@
To evaluate inductive abstraction and rule discovery, benchmarks like MM-IQ (Cai et al., 2025), LogicVista (Xiao et al., 2024), and PuzzleTest (Toh et al., 2025) present visual matrix-style puzzles inspired by Raven’s Progressive Matrices. These benchmark challenge models to infer abstract transformations, detect visual analogies, and fill in missing patterns by synthesizing structural invariants across multiple image frames. They emphasize high-level visual abstraction and the ability to generalize across non-linguistic reasoning patterns.
@@
A growing number of composite benchmarks have been proposed to evaluate generalized multimodal reasoning across varied tasks and domains. Representative datasets such as VRC-Bench (Thawakar et al., 2025) and MLLM-COMPBench (Kil et al., 2024) aggregate subtasks including visual QA, multi-hop reasoning, and tool-based inference, offering a unified setup for multi-skill evaluation. VCR-Bench (Qi et al., 2025) further extends this line into the video domain by incorporating annotated CoT steps and assessing both answer and reasoning quality across temporal tasks. Notably, M3CoT (Chen et al., 2024c) and MME-CoT (Jiang et al., 2025) include multimodal CoT annotations, enabling fine-grained process supervision during inference. Meanwhile, ZeroBench (Roberts et al., [n. d.]) targets zero-shot generalization across unseen tasks and modalities, testing model transferability and robustness without instruction tuning. Together, these benchmarks offer a comprehensive suite for evaluating holistic, structured, and cross-modal reasoning under diverse and complex inputs.
@@ 7.1. Limitations and Challenges
One of the most pressing issues is the sparsity and delayed nature of reward signals. To avoid reward hacking, most RL-enhanced MLLMs are trained using scalar final-task rewards—such as answer correctness or classification accuracy—that only reflect task outcomes but offer no intermediate feedback throughout the reasoning process. This leaves the model blind to structural errors in earlier reasoning steps. The most immediate consequence of this issue is overthinking, where the model generates excessively long or redundant reasoning paths, with intermediate steps potentially including irrelevant visual or textual cues. Although recent works (Zhang et al., 2025a; Deng et al., 2025b) have introduced process-oriented feedback through intermediate step verification or curriculum-style stage partitioning, these designs are often handcrafted, task-specific, and difficult to generalize across unseen modalities or new domains.
@@
Another limitation lies in the benchmark-centric and static nature of current evaluation paradigms. Many RL-based MLLMs are trained and tested on a limited range of benchmarks, which are typically narrow in scope and fail to reflect the complexity and unpredictability of real-world multimodal reasoning. As a result, models trained on such benchmarks, especially smaller-scale ones, tend to exhibit poor transferability when applied to dynamic environments or introduced to new modalities like audio or 3D spatial layouts.
@@
Beyond structural limitations, MLLM-RL pipelines face key challenges in optimization, multimodal alignment, and training infrastructure. A central issue is aligning diverse modalities such as images, text, audio, and spatial layouts under weak supervision. Real-world tasks often require complex cross-modal mappings with limited direct supervision, and designing reward functions to support consistent alignment, especially in open-ended scenarios, remains unsolved. Another challenge is modeling non-Markovian dependencies in reasoning trajectories. Unlike traditional RL, MLLM reasoning requires long-term consistency across modalities and steps. Although some work (Kang et al., 2025; Pan and Liu, 2025) has addressed this, optimization in non-Markovian spaces is prone to instability, noisy gradients, and unclear credit assignment. Finally, training-inference mismatch persists: models trained with fixed prompts and deterministic rewards are evaluated on unpredictable inputs, variable reasoning lengths, and ambiguous outputs—compromising real-world performance despite strong benchmarks.
@@ 7.2. Future Directions
To overcome these limitations and challenges, we propose several concrete directions for future research that can contribute to more robust, generalizable, and interactive RL in MLLMs:
@@
Future work should focus on designing multi-level reward signals that capture not only task outcomes but also the fidelity of reasoning and cross-modal coherence. For example, hybrid rewards can integrate accuracy, structure, and quality, thereby improving sample efficiency, interpretability, and policy stability.
@@
A promising research direction is the development of lightweight RL techniques suitable for resource-constrained models. This includes curriculum learning, KL-regularized off-policy methods, and contrastive reward estimation. Such strategies can facilitate broader adoption in scenarios where deploying large models is infeasible.
@@
Deploying MLLMs in spatially grounded environments such as robotics and AR/VR requires reinforcement learning strategies capable of handling physical constraints, causal relationships, and temporal dynamics. This includes integrating spatial consistency checks, collision detection, and user affordance modeling into the reward computation pipeline, as seen in MetaSpatial.